# Experimental Design for Active Transductive Inference in Large Language Models

### v1

@@ B.3 Proof of Theorem 5
Fix an $\mathbf{x}_{i}\in\mathcal{X}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$. Denote this new co-variance matrix as $\widehat{\mathbf{\Sigma}}_{t,i}$. Let $\widetilde{Y}_{t,i,j,1}=\mathbf{x}_{*}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,1}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t,i},\widehat{\mathbf{\Sigma}}_{t+1}\right)$ and $\epsilon_{t,i,j,1}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. This yields that $\widetilde{Y}_{t,i,j,1}\sim\mathcal{N}\left(\mathbf{x}_{*}^{\top}\widehat{\bm{\theta}}_{t+1},\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}\right)$.
Similarly $\widetilde{Y}_{t,i,j,2}=\mathbf{x}_{*}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,2}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t+1},\widehat{\mathbf{\Sigma}}_{t,i}\right)$ and $\epsilon_{t,i,j,2}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Therefore, we get that
$\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2}\sim\mathcal{N}(0,2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}).$
Now we proceed the same way as in Lemma 2 of Lalitha et al. (2023). Using Cochran’s theorem, we have that $\hat{\sigma}_{t,i}^{2}m/\sigma_{i}^{2}$ is a $\chi^{2}$ random variable with $m$ degrees of freedom. Then using (4.4) and Lemma 1 of Laurent & Massart (2000) we can show that
$\mathbb{P}\left(m-\frac{\hat{\sigma}_{s,t,i}^{2}m}{\sigma_{i}^{2}}\geq 2\sqrt{m\log(1/\delta)}\right)\leq\delta$ (6)
Dividing both sides of (6) in the probability by $m$, and multiplying by $\sigma_{i}^{2}$, we can get the following
$\mathbb{P}\left(\sigma_{i}^{2}(1-2\sqrt{\log(1/\delta)/m})\geq\hat{\sigma}_{s,t,i}^{2}\right)\leq\delta.$
Observe that $01-2\sqrt{\log(1/\delta)/m}>$ 0, we can divide both sides by it and get the first claim of the lemma. The second claim is proved by (4.3) in Laurent & Massart (2000), an immediate corollary of their Lemma 1, we have
$\mathbb{P}\left(\frac{\hat{\sigma}_{s,t,i}^{2}m}{\sigma_{i}^{2}}-m\geq 2\sqrt{m\log(1/\delta)}+2\log(1/\delta)\right)\leq\delta$
@@ Appendix C Additional Experiments and Results
We now briefly discuss the various datasets used in this work.
@@ C.1 Datasets
We now briefly describe the datasets used for our experiments. All the real-life datasets are from UCI (Markelle Kelly, 1988) and OpenML (Vanschoren et al., 2013) repositories. We use $4$ classification and $3$ regression datasets from UCI and OpenML. Additionally, we use $2$ custom datasets for movie names and entity names for classification task in our experiments. These are as follows:
@@
(1) Iris: We use this UCI dataset for classification task. This dataset consists of four features of flowers and three classes of flowers. We use all four features in the prompts as well as estimating the score for selecting the next action.
@@
(2) Banknote-authentication: We use this OpenML dataset for classification task. This dataset consists of five features of banknotes and two classes for identifying fake or original banknote. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action.
@@
(3) Balance-scale: We use this OpenML dataset for classification task. This dataset consists of five features of a scale and three classes of whether the scale tips left/right or is balanced. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action.
@@
(4) Thyroid-new: We use this OpenML dataset for classification task. This dataset consists of six features for thyroids and three classes. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action.
@@
(5) Movie-name: We use this custom dataset for classification task. This dataset consists of movie names across five genres (classes) romance, horror, thriller, sport, and action. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the movie names and the goal is to identify the common genre.
@@
(6) Movie-theme: We use this custom dataset for classification tasks in identifying a common theme between pairs of movies. This dataset consists of movie names across five themes (classes) good-vs-evil, man-vs-nature, redemption, Love conquers all, and coming-of-age. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the pair of movie names and the goal is to identify the common theme.
@@
(7) Entity-name: We use this custom dataset for classification task. This dataset consists of entity names across five entity types (classes) like mountains, seas, rivers, vehicles, and celebrities. Again, we convert the entity names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the entity names and the goal is to identify the entity type.
@@
(8) Fifa: We use this OpenML dataset for the regression task. This dataset consists of six features of players and the clubs they joined as targets. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action.
@@
(9) Machine-cpu: We use this OpenML dataset for regression tasks. This dataset consists of seven features of machine cpu and the target variable is the performance of the cpu. Out of these seven features, we use five features in the prompts as well as estimating the score for selecting the next action.
@@ C.1.1 ARC Experiment
In the recent works of Mirchandani et al. (2023); Srivastava et al. (2022) they showed that LLMs behave as general pattern matching machine. In fact they showed that LLMs can be used to solve tasks from Abstract Reasoning Corpus (ARC) tasks. In the following experiments we choose two such tasks: (1) ARC expansion and contraction experiment and (2) ARC rotation experiment.
(1) ARC expansion and contraction experiment: In the expansion and contraction experiment, there are two sets of input matrices of dimension $4\times 4$ which constitute the input feature space $\mathcal{X}$. The first set of input matrices are have integer values in their center $2\times 2$ cells while all the other cells are 00. The output space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ inner cells have moved to the $4$ corners. These matrices are termed as expansion matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the input feature space $\mathcal{X}$. Then the output space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells come to the center and all the other cell values are 00. These matrices are termed as contraction matrices. This is shown in Figure 5(a).
Therefore, the input feature space consist of both the expansion and contraction matrices. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 5(c).
(2) ARC rotation experiment: In the rotation experiment, there are again two sets of input matrices of dimension $4\times 4$ which constitute the input feature space $\mathcal{X}$. The first set of input matrices are have integer values in their four corner cells while all the other cells are 00. The output space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ corner cells values have moved $90^{\circ}$ in the clockwise direction. These matrices are termed as clockwise matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the input feature space $\mathcal{X}$. Then the output space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells have moved $15^{\circ}$ in the anti-clockwise direction and all the other cell values are 00. These matrices are termed as anit-clockwise matrices. This is shown in Figure 5(b).
Therefore, the input feature space consist of both the clockwise and anti-clockwise matrices. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 5(d).
@@
(c) Classification on Expansion and Contraction
(d) Classification on ARC Rotation Experiment
Figure 5: Experiment results with Falcon 40B on ARC tasks. For the classification tasks, we show the cumulative mis-classification error in the y-axis.
@@ C.1.2 PCFG Experiment
(1) PCFG add-subtract experiment: In the add-subtract experiment, there are two sets of input sequences of $4$ integers which constitute the input feature space $\mathcal{X}$. The first set of input sequences consist of odd integer values. The output space $\mathcal{Y}$ of this sequences of $4$ integers is a sequences of $5$ integers where the last integer is padded to the original sequence by adding one to the last odd integer. These sequences are termed as add sequences.
Similarly, the other set of input sequences consist of even integer values. The output space $\mathcal{Y}$ of this sequences of $4$ even integers is a sequences of $5$ integers where the last integer is padded to the original sequence by subtracting one from the last even integer. These sequences are termed as even sequences. This is shown in Figure 6(a).
Therefore, the input feature space consist of both the odd and even sequences. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 6(c).
(1) PCFG repeat experiment: In the repeat experiment, there are two sets of input sequences of $4$ integers which constitute the input feature space $\mathcal{X}$. The first set of input sequences consist of odd integer values. The output space $\mathcal{Y}$ of this sequences of $4$ integers is a sequences of $5$ integers where the last integer is padded to the original sequence by repeating the first odd integer. These sequences are termed as odd-repeat sequences.
Similarly, the other set of input sequences consist of even integer values. The output space $\mathcal{Y}$ of this sequences of $4$ even integers is a sequences of $5$ integers where the last integer is padded to the original sequence by repeating the second even integer. These sequences are termed as even-repeat sequences. This is shown in Figure 6(b).
Therefore, the input feature space consist of both the odd-repeat and even-repeat sequences. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 6(d).
@@
(c) Classification on PCFG add-subtract Experiment
(d) Classification on PCFG repeat Experiment
Figure 6: Experiment results with Falcon 40B on PCFG tasks. For the classification tasks, we show the cumulative mis-classification error in the y-axis.
C.2 Additional results
To provide a more quantitative comparison, we compute the mean and standard error of the cumulative error observed at the end of each trial run. As shown in Tables 1 and 2, GO and SAL are the absolute top 1 and 2 performing methods in 16 out of 20 tasks, and are close to the best-performing method on the rest 4 tasks.
Table 1: Cumulative error at the end of each run. The mean and standard errors for 20 repeat runs are shown for all classification datasets for 1-class and 2-class settings, with the top-1 method in bold. For movite-name, movie-theme, and entity-name, Instructor embedding is used as the feature for active experiment design.SAL and GO outperforms other baselines on the majority of the tasks.
Dataset iris banknote balance thyroid movie-name movie-theme entity-name
1-class classification
Uniform $9.05\pm 0.81$ $10.95\pm 0.45$ $12.40\pm 0.54$ $11.53\pm 0.80$ $12.05\pm 0.56$ $15.07\pm 0.57$ $13.40\pm 0.57$
Greedy-NN $13.43\pm 0.35$ $10.57\pm 0.29$ $12.75\pm 0.77$ $12.05\pm 0.53$ $13.18\pm 0.89$ $12.12\pm 1.17$ $15.07\pm 0.70$
D-opt $10.03\pm 0.76$ $10.22\pm 0.38$ $13.15\pm 0.63$ $12.93\pm 0.71$ $11.55\pm 0.54$ $15.53\pm 0.56$ $13.43\pm 0.48$
GO (ours) $7.12\pm 1.01$ $8.32\pm 0.71$ $9.78\pm 0.97$ $\mathbf{7.47\pm 0.80}$ $8.40\pm 0.54$ $\mathbf{11.20\pm 0.82}$ $\mathbf{9.32\pm 0.78}$
SAL (ours) $\mathbf{7.05\pm 0.93}$ $\mathbf{8.25\pm 0.74}$ $\mathbf{9.03\pm 0.91}$ $7.88\pm 0.43$ $\mathbf{8.25\pm 0.56}$ $11.55\pm 0.83$ $10.15\pm 0.76$
2-class classification
Uniform $\mathbf{8.79\pm 0.53}$ $10.57\pm 0.34$ $12.86\pm 0.43$ $12.19\pm 0.53$ $11.88\pm 0.57$ $14.51\pm 0.53$ $12.66\pm 0.40$
Greedy-NN $13.43\pm 0.27$ $10.72\pm 0.23$ $13.75\pm 0.39$ $12.41\pm 0.25$ $13.86\pm 0.63$ $\mathbf{13.80\pm 0.69}$ $14.35\pm 0.44$
D-opt $8.95\pm 0.52$ $10.51\pm 0.26$ $13.16\pm 0.35$ $11.47\pm 0.35$ $11.29\pm 0.44$ $14.55\pm 0.56$ $\mathbf{11.54\pm 0.46}$
GO (ours) $8.93\pm 0.50$ $10.80\pm 0.28$ $\mathbf{11.79\pm 0.25}$ $\mathbf{10.75\pm 0.45}$ $11.55\pm 0.48$ $14.06\pm 0.39$ $11.69\pm 0.35$
SAL (ours) $8.93\pm 0.49$ $\mathbf{10.40\pm 0.18}$ $12.75\pm 0.33$ $11.12\pm 0.37$ $\mathbf{11.18\pm 0.68}$ $14.25\pm 0.36$ $12.29\pm 0.49$
Table 2: Cumulative error at the end of each run. Mean and standard error for 20 repeat runs are shown for all ARC, PCFG and regression datasets, with top-1 method in bold. SAL and GO outperform other baselines on the majority of the tasks.
Dataset arc_expand_contract arc_rotation pcfg_add_subtract pcfg_repeat fifa CPU
Uniform $3.60\pm 0.33$ $4.25\pm 0.19$ $14.25\pm 0.70$ $\mathbf{6.25\pm 0.83}$ $31.14e4\pm 4.36e$ 3 $33.67e4\pm 3.61e$ 3
Greedy-NN $2.55\pm 0.38$ $5.00\pm 0.00$ $19.85\pm 0.08$ $18.65\pm 0.31$ $42.39e4\pm 1.15e$ 4 $41.08e5\pm 1.37e$ 4
D-opt $3.90\pm 0.27$ $4.15\pm 0.21$ $14.45\pm 0.78$ $9.90\pm 0.74$ $32.74e4\pm 3.36e$ 3 $32.99e4\pm 2.62e$ 3
GO (ours) $1.50\pm 0.26$ $2.35\pm 0.30$ $\mathbf{10.70\pm 0.68}$ $\mathbf{6.25\pm 0.72}$ $\mathbf{6.78e3\pm 2.07e3}$ $\mathbf{1.75e4\pm 2.05e3}$
SAL (ours) $\mathbf{1.00\pm 0.21}$ $\mathbf{1.85\pm 0.18}$ $11.35\pm 0.93$ $6.90\pm 0.61$ $38.80e3\pm 2.99e$ 3 $31.77e4\pm 2.16e$ 3

### v2

@@ B.4 Proof of Theorem 2
Fix an $\mathbf{x}_{i}\in\mathcal{X}_{\text{examples}}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$. Denote this new co-variance matrix as $\widehat{\mathbf{\Sigma}}_{t,i}$ such that
$\widehat{\mathbf{\Sigma}}_{t,i}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)\right)^{-1}.$
Let $\widetilde{Y}^{(1)}_{t,i,j}=\mathbf{x}_{*,k}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,1}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t},\widehat{\mathbf{\Sigma}}_{t+1}\right)$ and $\epsilon_{t,i,j,1}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. This yields that $\widetilde{Y}^{(1)}_{t,i,j}\sim\mathcal{N}\left(\mathbf{x}_{*,k}^{\top}\widehat{\bm{\theta}}_{t+1},\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}\right)$. Similarly $\widetilde{Y}^{(2)}_{t,i,j}=\mathbf{x}_{*,k}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,2}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t+1},\widehat{\mathbf{\Sigma}}_{t,i}\right)$ and $\epsilon_{t,i,j,2}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Therefore, we get that
$\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k}\sim\mathcal{N}(0,2\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}).$
Now we proceed the same way as in Lemma 2 of Lalitha et al. [2023]. Using Cochran’s theorem, we have that $\hat{\sigma}_{t,i,k}^{2}m/\sigma_{i,k}^{2}$ is a $\chi^{2}$ random variable with $m$ degrees of freedom. Then using (4.4) and Lemma 1 of Laurent and Massart [2000] we can show that
$\mathbb{P}\left(m-\frac{\hat{\sigma}_{t,i,k}^{2}m}{\sigma_{i,k}^{2}}\geq 2\sqrt{m\log(1/\delta)}\right)\leq\delta$ (5)
Dividing both sides of (5) in the probability by $m$, and multiplying by $\sigma_{i,k}^{2}$, we can get the following
$\mathbb{P}\left(\sigma_{i,k}^{2}(1-2\sqrt{\log(1/\delta)/m})\geq\hat{\sigma}_{t,i,k}^{2}\right)\leq\delta.$
Observe that $01-2\sqrt{\log(1/\delta)/m}>$ 0, we can divide both sides by it and get the first claim of the lemma. The second claim is proved by (4.3) in Laurent and Massart [2000], an immediate corollary of their Lemma 1, we have
$\mathbb{P}\left(\frac{\hat{\sigma}_{t,i,k}^{2}m}{\sigma_{i,k}^{2}}-m\geq 2\sqrt{m\log(1/\delta)}+2\log(1/\delta)\right)\leq\delta$
@@ Appendix C Additional Experiments and Results
We use NVIDIA GeForce RTX 3090 GPU with 24GB RAM to load the Large Language Models for inference. The Mistral-7B model requires less than 16GB RAM, and Vinuna-13B model requires less than 22 GB RAM during execution. To run Falcon-40B model we use AWS ml-g5.12xlarge machine. To run the full set of experiments it takes 24-27 hours of compute job. We now briefly discuss the various datasets used in this work.
@@ C.1 Datasets
We now briefly describe the datasets used for our experiments. All the real-life datasets are from UCI [Markelle Kelly, 1988] and OpenML [Vanschoren et al., 2013] repositories. We use $4$ classification and $3$ regression datasets from UCI and OpenML. Additionally, we use $2$ custom datasets for movie names and entity names for classification task in our experiments. These are as follows:
@@
(1) Iris: We use this UCI dataset for classification task. This dataset consists of four features of flowers and three classes of flowers. We use all four features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $150$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
@@
(2) Banknote-authentication: We use this OpenML dataset for classification task. This dataset consists of five features of banknotes and two classes for identifying fake or original banknote. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $150$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
@@
(3) Balance-scale: We use this OpenML dataset for classification task. This dataset consists of five features of a scale and three classes of whether the scale tips left/right or is balanced. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $625$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
@@
(4) Thyroid-new: We use this OpenML dataset for classification task. This dataset consists of six features for thyroids and three classes. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $215$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
@@
(5) Movie-name: We use this custom dataset for classification task. This dataset consists of movie names across five genres (classes) romance, horror, thriller, sport, and action. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the movie names and the goal is to identify the common genre. The dataset consists of $100$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
@@
(6) Movie-theme: We use this custom dataset for classification tasks in identifying a common theme between pairs of movies. This dataset consists of movie names across five themes (classes) good-vs-evil, man-vs-nature, redemption, Love conquers all, and coming-of-age. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the pair of movie names and the goal is to identify the common theme. The dataset consists of $100$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
@@
(7) Entity-name: We use this custom dataset for classification task. This dataset consists of entity names across five entity types (classes) like mountains, seas, rivers, vehicles, and celebrities. Again, we convert the entity names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the entity names and the goal is to identify the entity type. The dataset consists of $100$ pairs of instances. We randomly choose $K=20$ pairs as test examples and the remaining instances as training examples.
@@
(8) Fifa: We use this OpenML dataset for the regression task. This dataset consists of six features of players and the clubs they joined as targets. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $18063$ instances. We randomly choose $K=20$ test examples and another $200$ examples as training examples.
@@
(9) Machine-cpu: We use this OpenML dataset for regression tasks. This dataset consists of seven features of machine cpu and the target variable is the performance of the cpu. Out of these seven features, we use five features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $209$ instances. We randomly choose $K=20$ test examples and the remaining examples as training examples.
@@ C.1.1 ARC Experiment
In the recent works of Mirchandani et al. [2023], Srivastava et al. [2022] they showed that LLMs behave as general pattern-matching machine. In fact they showed that LLMs can be used to solve tasks from Abstract Reasoning Corpus (ARC) tasks. In the following experiments, we choose two such tasks: (1) ARC expansion and contraction experiment and (2) ARC rotation experiment.
(1) ARC expansion and contraction experiment: In the expansion and contraction experiment, there are two sets of matrices of dimension $4\times 4$ which constitute half the examples of the feature space $\mathcal{X}$. The first set of input matrices have integer values in their center $2\times 2$ cells while all the other cells are 00. The label space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ inner cells have moved to the $4$ corners. These matrices are termed as expansion matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the remaining examples in $\mathcal{X}$. Then the label space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells come to the center and all the other cell values are 00. These matrices are termed as contraction matrices. This is shown in Figure 1(a).
Therefore, the feature space $\mathcal{X}$ consists of both the expansion and contraction matrices. At every trial, and $n$ training examples and $K$ test examples are chosen randomly from $\mathcal{X}$. Then we run all baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
(2) ARC rotation experiment: In the rotation experiment, there are again two sets of matrices of dimension $4\times 4$ which constitute half the examples of the feature space $\mathcal{X}$. The first set of matrices are have integer values in their four corner cells while all the other cells are 00. The label space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ corner cell values have moved $90^{\circ}$ in the clockwise direction. These matrices are termed as clockwise matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the remaining examples in the feature space $\mathcal{X}$. Then the label space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells have moved $15^{\circ}$ in the anti-clockwise direction and all the other cell values are 00. These matrices are termed as anti-clockwise matrices. This is shown in Figure 1(b).
Therefore, the feature space $\mathcal{X}$ consists of both the clockwise and anti-clockwise matrices. At every trial, $n$ training examples and $K$ test examples are chosen randomly. Then we run all the baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
@@
Figure 1: Explanation of ARC tasks
@@ C.1.2 PCFG Experiment
(1) PCFG add-subtract experiment: In the add-subtract experiment, there are two sets of sequence of $4$ integers. The first set of sequence of $4$ integers consists of odd integer values which constitute half the examples in the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of this sequence of $4$ odd examples is sequences of $5$ integers where the last integer is padded to the original sequence by adding one to the last odd integer. These sequences are termed as add examples.
Similarly, the other set of examples consists of a sequence of $4$ even integer values which constitute the remaining examples in the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of this sequence of $4$ even integer examples is a sequence of $5$ integers where the last integer is padded to the original sequence by subtracting one from the last even integer. These examples are termed as even examples. This is shown in Figure 2(a).
Therefore, the feature space $\mathcal{X}$ consists of both the odd and even sequence of $4$ integer value examples. At every trial, $n$ training and $K$ test examples are chosen randomly. Then we run all the baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
(2) PCFG repeat experiment: In the repeat experiment, there are two sets of sequence of $4$ integers. The first set of sequence of $4$ odd integer values constitute half the examples of the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of this sequence of $4$ integers examples is a sequence of $5$ integers where the last integer is padded to the original sequence by repeating the first odd integer. These sequences are termed as odd-repeat examples.
Similarly, the other set of examples consists of sequence of $4$ even integer values which constitute the remaining examples in the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of these sequence of $4$ integer value examples is a sequence of $5$ integers where the last integer is padded to the original sequence by repeating the second even integer. These examples are termed as even-repeat examples. This is shown in Figure 2(b).
Therefore, the feature space $\mathcal{X}$ consist of both the odd-repeat and even-repeat examples. At every trial, $n$ training examples and $K$ test examples are chosen randomly. Then we run all the baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
@@
Figure 2: Explanation of PCFG task.